TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-04-21 23:24:00
subject: 4\08 Pt 1 ISS On Orbit Status 08-04-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

09 Apr 2003

ISS On-Orbit Status 8 Apr 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.

After wakeup at 2:00am (all times EDT), morning inspection and
breakfast, CDR Ken Bowersox and FE-2/SO Don Pettit began final
preparations for today's EVA, the second Stage EVA of their
increment (the first was on 1/15/03). After an initial 80-min.
prebreathe by the spacewalkers with PHA (pre-breathe hose assembly)
masks for O2 (oxygen) plus exercise on the workload-controlled CEVIS
for denitrogenation of their bodies, the crew ingressed the
Airlock's (A/L) Crewlock compartment at about 4:30am EDT.  The actual
EVA began four hours later, at 8:40am.   [In the A/L, Budarin closed
the node/Node hatch and assisted Sox and Don with EMU (extravehicular
mobility unit) donning.  At 5:00am, Nikolai started the first
depress, from 14.7 psi (760 mmHg) to 10.2 (527), to support the
subsequent EMU purge for eliminating any remaining nitrogen from the
suits.  Air pressure was then briefly returned to 14.7 psi for
Budarin to exit the A/L, followed by an additional 60-min. in-suit
prebreathe period.  Finally, the A/L was depressed in stages (5 psi,
then 3.1 psi) to zero, with hatch opening at 8:40am.  The 60-min.
in-suit prebreathe includes 20 min. of safety margin; it can be
reduced to 40 min. in the future when an operational ISD (in-suit
Doppler) instrument is available to actually help measure VGE (venous
gas emboli) and validate the original prebreathe protocol.]

All EVA objectives were successfully accomplished.  [Sox reconfigured
electrical connectors at the S0/S1 truss BBC (bus bolt controller)
panel to provide additional inhibits preventing inadvertent SSAS
(segment-to-segment attachment system) release, while Don removed and
replaced the failed RPCM (remote power controller module) 3A, one of
two for the MT (mobile transporter) cart, restoring its power
redundancy.  Next task was reconfiguring CMG (control moment
gyroscope) connections at the Z1 truss "rat's nest", port and
starboard, to ensure that no single EPS (electrical power system)
failure will result in loss of two of the three operating CMGs,
thereby crippling the U.S. attitude control system (ACS).  The
spacewalkers then installed two remedial SPD (spool position device)
clamps at the Lab heat exchanger Loop-A QD (quick disconnect) ammonia
lines, addressing generic hydraulic lockup concerns with wet fluid
QDs.  They also successfully freed the light stanchion of the S1
CETA (crew equipment translation aid) railcart, which had remained
stuck after their aborted light installation during the first Stage
EVA, by tapping it with a Russian hammer, then attached the CETA
light fixture to the pole.  When all primary objectives were
accomplished after 4h 40 min., Sox and Don performed a number of
extra "get-ahead" tasks, such as relocation of tools, clamps and
tethers from the exterior of the Node and Z1 truss to toolboxes on
the A/L, and taking photographs of selected outside views.  After EVA
cleanup, the crew returned to the A/L, completing the spacewalk at
3:07pm, after a total duration of 6h 27m.  It was the 51st EVA of
station assembly, the 26th from the ISS (the others from Shuttle),
and the 17th from the Joint A/L (the others from the Russian DC-1),
but only the third with no Shuttle present. Total US EVA time for ISS
is now 318h 38m (151h 23m from the A/L).] 

In support of the spacewalk, during Daily Orbit 14 (DO14) and DO6 ISS
flight control went to RS (Russian segment) motion control and SM
thrusters twice for momentum-desaturation of CMG-1 and -2,- once
before, the other after the EVA. The next attitude control handover
to RS will be on Thursday, 4/10, for the second Progress reboost
(6:34am EDT).

Prior to the EVA, at 8:30am, MCC-H commanded CMG-1 and -2 shut-down
in preparation for the CMG power reconfiguration at the Z1 truss by
Sox and Don.  [The gyros were not braked to zero but remained in
unpowered "coast" mode, with RPC (remote power controller) inhibits,
slowing from the original 6600 RPM to 5000 RPM for quick readiness in
case of a contingency.  They were powered up and spun up again at
about 10:40am.] 

Last night's swap-out of the two GNC MDMs (guidance, navigation &
control computers) was completed without incident.  [The new GNC
configuration table in support of the CMG RPCM redesignation during
the EVA was successfully loaded in both MDMs.  Afterwards, uplinked
contingency measures by the crew for a primary GNC MDM failure were
no longer needed.  GNC-1 is now primary again.]

The thermal control system (TCS) was successfully configured to
support the SPD installation on the external Loop A heat exchanger
lines. 

To minimize the spacewalkers' required post-EVA time on the schedule,
METOX (metal oxide) adsorbent canister regeneration and EMU water
recharge was deferred to tomorrow, 4/9.

Budarin, who "presided" over the EVA as IV (intravehicular
crewmember) from the inside, changed out VTR tapes for recording
SSRMS video imagery and reconfigured the appropriate communications
modes (2B and 1A) before and after the excursion.

Nikolai also performed the regular daily maintenance of the SOZh life
support system in the SM.  Early in the morning, he conducted his
regular quick inspection of the BIO-5 Rasteniya-2/Lada-2 ("Plants-2")
plant growth experiment.

Tomorrow the crew will have a half-day off for rest.

Launch of Soyuz TMA-2/6S on 4/26 is currently scheduled for 9:53am
local Baikonur time (7:53am Moscow time, 11:53pm EDT on 4/25).  6S
docking is expected at 4/28, 1:54am EDT (9:54am Moscow time).

Today's CEO (crew earth observations) targets, currently restricted
by flight rule constraints limiting the use of the science window in
the Lab, were Ganges River Delta (good ISS pass here for a near-nadir
mapping passage over this large, complex delta feature), Dhaka,
Bangladesh (an excellent nadir pass over the Bangladeshi capital
city), Johannesburg, South Africa (South Africa's major industrial
city lied just left of track), Tunis, Tunisia (the Tunisian capital
is located on the coast, just left of track on this pass), Lisbon,
Portugal (this was a nadir pass over Portugal's capital, located on a
large estuary), and Lima, Peru (the Peruvian capital is located on
the coast.  This nadir pass was early enough in the day to precede
daytime cloud formation). CEO images can be viewed at the websites
http://eol.jsc.nasa.gov and http://earthobservatory.nasa.gov

U.S. and Russian Segment Status (as of 1:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS):

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.